# Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework

### v1

@@ Abstract
It was implemented within many different frameworks, some using non-relativistic (e.g. Refs. [3, 4] ) or relativistic (e.g. Refs. [5, 6]) density functionals, other exploiting the Green’s function formalism (e.g. Refs. [7, 8]). The latter approach allows one to go beyond the harmonic approximation by coupling the RPA modes to complex configurations. Other extensions using Skyrme forces were proposed [9]. The investigations based on modern realistic potentials are few and rather recent [10, 11, 12, 13, 14, 15].
@@
Most of the recipes for obviating this shortcoming were enumerated and developed long ago by Rowe [16, 17]. They consist in reintroducing the correlations into the ground state without spoiling the simple structure of the RPA eigenvalue equations. Few investigations were conducted along this line since then [18, 19, 20]. An important extension was achieved and applied to metal clusters in Refs. [21, 22].
It is worth mentioning a more recent approach using the realistic Argonne v18 potential [23], as well as a dressed (D) RPA study performed within a self-consistent Green’s function scheme [14]. A variety of works aiming to go beyond RPA is presented and discussed in a recent review [24].
Other approaches for computing ground-state correlations are available. We mention an extension of TDA [25], a time-dependent density matrix (TDDM) formalism [26], a many-body perturbative approach [27], in-medium similarity renormalization group (IMSRG) [28] and an equation of motion multiphonon method (EMPM) [29].
@@
In the present work, we perform a systematic study of bulk properties, nuclear responses, and spectra of a large number of nuclei ranging from 4He to 208Pb. We use the chiral potential N2LOGO(394) [30], incorporating the three-body force, to generate a self-consistent HF basis and then solve the eigenvalue equations within TDA, RPA, and RRPA.
To our knowledge, this is the first ab initio extension of RPA which removes the instabilities induced by the QBA. It accounts for the nuclear bulk properties and provides a unified scheme where energy levels and responses are referred to the true correlated (rather than unperturbed) ground state.
@@
Note that in the limit the standard RPA expressions are recovered.
@@
The systematics presented in Fig. 2 prove that HF underestimates severely the binding energy of all studied nuclei. On the other hand, a strong over-binding is obtained once we add the contribution from the RPA ground-state correlations. The quenching action of the RRPA restores the consistency with the empirical values. The close agreement of the RRPA energies with the corresponding quantities obtained in HF + second-order perturbation theory (PT) is noteworthy.
RPA tends to overestimate also the nuclear radii (Fig. 2). RPA restores the agreement between theoretical and empirical quantities.
Next, we discuss the effect of renormalisation on selected excited states. In RPA, few levels of the 16O and 90Zr spectra are strongly pushed down toward the ground state (Fig. 3). This is a general feature occurring in all nuclei investigated. In some of them, like 40Ca, the lowest level collapses and the energy becomes imaginary. These instabilities are removed within the RRPA and the consistency with the experiments is restored.
Further insight can be obtained from the transition probabilities. The ground-state correlations, accounted for in RRPA, do not produce any significant further fragmentation of the strength with respect to RPA. On the other hand, because of the energy downshift and damping of the main E1 peak they cause, the energy-weighted sum rule, preserved in RPA, is sensibly underestimated (Fig. 4). This impact is to be attributed to the depletion of few single particle (hole) states since the majority of the states involved in the transitions, being far from the Fermi surface, are empty (full) (Fig.5).
The depletion of single-particle levels around such a surface affects strongly the low-lying octupole transitions (Table 1). In fact, the ground-state reduced strength of the transition to the 3, practically negligible in TDA, increases dramatically as we move to RPA which systematically overestimates the measured values. These are sensibly underestimated in RRPA as a consequence of the drastic quenching of the amplitudes (Fig. 5).
@@
We can therefore conclude that, within the limits of its validity, the present self-consistent RRPA approach offers, from first principles, a simple reliable unified systematic of bulk and spectroscopic properties of finite closed (sub-)shell nuclei, and is especially useful in the regions of medium and heavy nuclei, not easily accessible to more complete ab initio approaches.
@@ Acknowledgements.
This work is supported by the Czech Science Foundation (Czech Republic), P203-23-06439S. Computational resources were provided by the CESNET LM2015042 and the CERIT Scientific Cloud LM2015085, under the program ”Projects of Large Research, Development, and Innovations Infrastructures”.References
- Rowe [2010] D. J. Rowe, Nuclear collective motion: models and theory (World Scientific, 2010).
- Ring and Schuck [2004] P. Ring and P. Schuck, The nuclear many-body problem (Springer Science & Business Media, 2004).
- Bender et al. [2003] M. Bender, P.-H. Heenen, and P.-G. Reinhard, Rev. Mod. Phys. 75, 121 (2003).
- Roca-Maza and Paar [2018] X. Roca-Maza and N. Paar, Progress in Particle and Nuclear Physics 101, 96 (2018).
- Vretenar et al. [2005] D. Vretenar, A. Afanasjev, G. Lalazissis, and P. Ring, Physics Reports 409, 101 (2005).
- Paar et al. [2003] N. Paar, P. Ring, T. Nikšić, and D. Vretenar, Phys. Rev. C 67, 034312 (2003).
- Lenske and Wambach [1990] H. Lenske and J. Wambach, Physics Letters B 249, 377 (1990).
- Litvinova [2023] E. Litvinova, The European Physical Journal A 59, 291 (2023).
- Gambacurta et al. [2015] D. Gambacurta, M. Grasso, and J. Engel, Phys. Rev. C 92, 034303 (2015).
- Paar et al. [2006] N. Paar, P. Papakonstantinou, H. Hergert, and R. Roth, Phys. Rev. C 74, 014318 (2006).
- Papakonstantinou and Roth [2010] P. Papakonstantinou and R. Roth, Phys. Rev. C 81, 024317 (2010).
- Hergert et al. [2011] H. Hergert, P. Papakonstantinou, and R. Roth, Phys. Rev. C 83, 064317 (2011).
- Wu et al. [2018] Q. Wu, B. S. Hu, F. R. Xu, Y. Z. Ma, S. J. Dai, Z. H. Sun, and G. R. Jansen, Phys. Rev. C 97, 054306 (2018).
- Raimondi and Barbieri [2019] F. Raimondi and C. Barbieri, Phys. Rev. C 99, 054327 (2019).
- Beaujeault-Taudière et al. [2023] Y. Beaujeault-Taudière, M. Frosini, J.-P. Ebran, T. Duguet, R. Roth, and V. Somà, Phys. Rev. C 107, L021302 (2023).
- Rowe [1968a] D. J. Rowe, Phys. Rev. 175, 1283 (1968a).
- Rowe [1968b] D. J. Rowe, Rev. Mod. Phys. 40, 153 (1968b).
- Ellis [1987] P. Ellis, Nuclear Physics A 467, 173 (1987).
- Karadjov et al. [1993] D. Karadjov, V. Voronov, and F. Catara, Physics Letters B 306, 197 (1993).
- Catara et al. [1994] F. Catara, N. Dinh Dang, and M. Sambataro, Nuclear Physics A 579, 1 (1994).
- Catara et al. [1996] F. Catara, G. Piccitto, M. Sambataro, and N. Van Giai, Phys. Rev. B 54, 17536 (1996).
- Catara et al. [1998] F. Catara, M. Grasso, G. Piccitto, and M. Sambataro, Phys. Rev. B 58, 16070 (1998).
- Roth and Paar [2007] P. P. R. Roth and N. Paar, Physical Review C 75, 014310 (2007).
- Schuck et al. [2021] P. Schuck, D. S. Delion, J. Dukelsky, M. Jemai, E. Litvinova, G. Röpke, and M. Tohyama, Physics Reports 929, 1 (2021).
- Minato [2016] F. Minato, Phys. Rev. C 93, 044319 (2016).
- Barton et al. [2021] M. Barton, P. Stevenson, and A. Rios, Phys. Rev. C 103, 064304 (2021).
- Hu et al. [2016] B. S. Hu, F. R. Xu, Z. H. Sun, J. P. Vary, and T. Li, Phys. Rev. C 94, 014303 (2016).
- Hergert et al. [2016] H. Hergert, S. Bogner, T. Moris, A. Schwenk, and K. Tsukiyama, Physics Reports 621, 165 (2016), ISSN 0370-1573.
- De Gregorio et al. [2017a] G. De Gregorio, J. Herko, F. Knapp, N. Lo Iudice, and P. Veselý, Phys. Rev. C 95, 024306 (2017a).
- Jiang et al. [2020] W. G. Jiang, A. Ekström, C. Forssén, G. Hagen, G. R. Jansen, and T. Papenbrock, Phys. Rev. C 102, 054301 (2020).
- Miyagi [2023] T. Miyagi, The European Physical Journal A - Hadrons and Nuclei 59, 150 (2023).
- Miyagi et al. [2022] T. Miyagi, S. R. Stroberg, P. Navrátil, K. Hebeler, and J. D. Holt, Phys. Rev. C 105, 014302 (2022).
- De Gregorio et al. [2021] G. De Gregorio, F. Knapp, N. Lo Iudice, and P. Veselý, Physics Letters B 821, 136636 (2021).
- De Gregorio et al. [2017b] G. De Gregorio, J. Herko, F. Knapp, N. Lo Iudice, and P. Veselý, Phys. Rev. C 95, 024306 (2017b).
- [35] National nuclear data center database, URL https://www.nndc.bnl.gov/nudat3/.
- Angeli and Marinova [2013] I. Angeli and K. Marinova, Atomic Data and Nuclear Data Tables 99, 69 (2013), ISSN 0092-640X.
- Goriely et al. [2019] S. Goriely, P. Dimitriou, M. Wiedeking, T. Belgya, R. Firestone, J. Kopecky, M. Krtička, V. Plujko, R. Schwengner, S. Siem, et al., The European Physical Journal A 55 (2019), ISSN 1434-601X.
- Spear [1989] R. H. Spear, Atomic Data and Nuclear Data Tables 42, 55 (1989), ISSN 0092-640X.
- Barrett et al. [2013] B. R. Barrett, P. Navrátil, and J. P. Vary, Progress in Particle and Nuclear Physics 69, 131 (2013).
- Hagen et al. [2014] G. Hagen, T. Papenbrock, M. Hjorth-Jensen, and D. J. Dean, Rep. Prog. Phys. 77, 096302 (2014).
- De Gregorio et al. [2016] G. De Gregorio, F. Knapp, N. Lo Iudice, and P. Veselý, Phys Rev. C 93, 044314 (2016).

### v2

@@ Abstract
It is an essential tool for investigating nuclear spectra and response functions, including pygmy and giant resonances, as well as weak-interaction processes that are of significant importance in astrophysical contexts.
RPA was implemented within many different frameworks, some using non-relativistic (e.g. Refs. [5, 36]) or relativistic (e.g. Refs. [43, 31]) density functionals, other exploiting the Green’s function formalism (e.g. Refs. [24, 25]). The latter approach allows one to go beyond the harmonic approximation by coupling the RPA modes to complex configurations. Other extensions using Skyrme forces were proposed [16]. The investigations based on modern realistic potentials are few and rather recent [30, 33, 20, 44, 34, 4].
@@
Most of the recipes for obviating this shortcoming were enumerated and developed long ago by Rowe [38, 37]. A more updated list of papers devoted to this task is presented and discussed in a recent review [41]. They consist in reintroducing the correlations into the ground state without spoiling the simple structure of the RPA eigenvalue equations. Few investigations were conducted along this line since then [15, 23, 7]. An important extension was achieved and applied to metal clusters in Refs. [9, 8].
The method so reformulated was applied to nuclei by Papakonstantiou et al. [32] using the UCOM potential derived from the Argonne V18 nucleon-nucleon (NN) interaction through a unitary transformation. The effect of the ground-state correlations on several multipole responses resulted to be modest.
Several alternative approaches accounting for ground-state correlations are available. We mention an extension of TDA [26], a time-dependent density matrix (TDDM) formalism [3], a many-body perturbative approach [21], coupled-cluster(CC) [18], in-medium similarity renormalization group (IMSRG) [19], and an equation of motion multiphonon method (EMPM) [11].
@@
In the present work, we adopted the method developed in Ref. [8] and extended to nuclei in Ref. [32] to carry out a systematic study of bulk properties, nuclear responses, and spectra of a large number of closed (sub-)shell nuclei ranging from 4He to 208Pb.
We use the chiral potential N2LOGO(394) [22], incorporating the three-body force, to generate a self-consistent HF basis and then solve the eigenvalue equations within TDA, RPA, and RRPA.
To our knowledge, this is the first ab initio renormalized version of RPA that directly employs the state-of-art two- and three-body interactions without intermediate smoothing steps. As we shall show, it describes within a unified framework nuclear bulk properties with accuracy comparable to the one achieved by the CC and IMSRG methods and yields, within the limits imposed by 1p-1h space, complete spectra and responses referred to the correlated ground state. Moreover, it covers all sub-shell-closed nuclei across the nuclear chart, including the heavy systems.
All these results were obtained at a computational cost significantly lower than the one spent by the available IMSRG solver [40].
Thus, the present approach offers a new efficient manageable tool for a systematic study of low and high energy nuclear properties from first principles. It should have, therefore, an important impact on nuclear structure and nuclear astrophysics.
@@
In the limit , the standard RPA formula is recovered.
@@
The systematics presented in Fig. 2 prove that HF underestimates severely the binding energy of all studied nuclei. Second-order perturbation theory (PT) refines the calculated energies, significantly reducing discrepancies with the experimental values. On the other hand, a strong over-binding is obtained once we add the contribution from the RPA ground-state correlations. The quenching action of the RRPA restores the consistency with the empirical values.
A more detailed analysis (Table 1) shows that RRPA reproduces the binding energies with an overall accuracy comparable to the one achieved within the CC and IMSRG approaches.
| RRPA | PT | IMSRG(2) | CC | Exp. | |
| 124.7 | 124.4 | 126.2 | 127.5 | 127.62 | |
| 168 | 164 | 167 | 169 | 168.96 | |
| 338 | 338 | 339 | 346 | 342.05 | |
| 416 | 410 | 415 | 420 | 416.00 | |
| 634 | 624 | 629 | 639 | 641.55 | |
| 773 | 764 | 768 | 782 | 783.90 | |
| 807 | 793 | 804 | 818 | 825.30 | |
| 1045 | 1021 | 1026 | 1043 | 1102.84 |
| RRPA | CC | Exp. | |
| 6.3 | 5.6 | 6.13 | |
| 3.8 | 3.9 | 4.79 | |
| 4.6 | 4.1 | 3.83 |
RRPA restores also the agreement between theoretical and empirical radii overestimated in RPA (Fig. 2).
Next, we discuss the effect of the renormalization on selected excited states. In RPA, few levels of the 16O and 90Zr spectra are strongly pushed down toward the ground state (Fig. 3). This is a general feature of all investigated nuclei. In some of them, like 40Ca, the lowest level collapses and the energy becomes imaginary. These instabilities are removed within the RRPA and the consistency with the experiments is restored.
As shown in Table 2, the low lying levels, obtained using the same potential and model space, are reproduced with comparable accuracy in RRPA and CC [22].
The RRPA monopole energy centroids are also comparable to the corresponding quantities obtained within recent LIT IMSRG/CC approaches [6].
Further insight can be obtained from the transition probabilities. The ground-state correlations, accounted for in RRPA. do not further produce any significant fragmentation of the strength with respect to RPA. On the other hand, because of the energy downshift and damping of the main E1 peak they cause, the energy-weighted sum rule, preserved in RPA, is sensibly underestimated (Fig. 4). This impact is to be attributed to the depletion of few single particle (hole) states since the majority of the states involved in the transitions, being far from the Fermi surface, are empty (full) (Fig.5).
The depletion of single-particle levels around such a surface strongly affects the low-lying octupole transitions (Table 3). In fact, the ground-state reduced strength of the transition to the 3, practically negligible in TDA, increases dramatically as we move to RPA which systematically overestimates the measured values. These are sensibly underestimated in RRPA as a consequence of the drastic quenching of the amplitudes (Fig. 5).
@@
We can therefore conclude that, within the limits of its validity, the present self-consistent RRPA approach offers, from first principles, a simple reliable unified systematic of bulk and spectroscopic properties of finite closed (sub-)shell nuclei at a moderate numerical computational cost, and is especially useful in the regions of medium and heavy nuclei, not easily accessible to the other well established ab initio approaches. Hopefully, it might open a new route, based on the direct use of bar modern potentials, to the large community adopting mean field approaches.
@@ Acknowledgements.
This work is supported by the Czech Science Foundation (Czech Republic), P203-26-21972S. Computational resources were provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic, and by the ELIXIR-CZ project (ID:90255).References
- [1] (2013) Table of experimental nuclear ground state charge radii: an update. Atomic Data and Nuclear Data Tables 99 (1), pp. 69 – 95. External Links: Document, ISSN 0092-640X Cited by: Figure 1, Figure 2.
- [2] (2013) Ab initio no core shell model. Progress in Particle and Nuclear Physics 69, pp. 131. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [3] (2021-06) Nuclear ground states in a consistent implementation of the time-dependent density matrix approach. Phys. Rev. C 103, pp. 064304. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [4] (2023-02) Zero- and finite-temperature electromagnetic strength distributions in closed- and open-shell nuclei from first principles. Phys. Rev. C 107, pp. L021302. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [5] (2003-01) Self-consistent mean-field models for nuclear structure. Rev. Mod. Phys. 75, pp. 121–180. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [6] (2025) Ab initio calculations of monopole sum rules: from finite nuclei to infinite nuclear matter. arXiv:2511.15836. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [7] (1994) Ground-state correlations beyond rpa. Nuclear Physics A 579 (1), pp. 1–12. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [8] (1998-12) Self-consistent determination of the one-body density matrix and particle-hole excitations. Phys. Rev. B 58, pp. 16070–16075. External Links: Document Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [9] (1996) Towards a self-consistent random-phase approximation for fermi systems. Phys. Rev. B 54 (24), pp. 17536–17546. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [10] Data extracted using the NNDC On-line Data Service from the ENSDF database. Cited by: Table 1.
- [11] (2017-02) Ground-state correlations within a nonperturbative approach. Phys. Rev. C 95, pp. 024306. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [12] (2017-02) Ground-state correlations within a nonperturbative approach. Phys. Rev. C 95, pp. 024306. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [13] (2021) Removal of the center of mass in nuclei and its effects on 4he. Physics Letters B 821, pp. 136636. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [14] (2016) Self-consistent quasiparticle formulation of a multiphonon method and its application to the neutron-rich nucleus. Phys Rev. C 93, pp. 044314. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [15] (1987) Particle-hole and particle-particle rpa ground state correlations. Nuclear Physics A 467 (2), pp. 173–184. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [16] (2015-09) Subtraction method in the second random-phase approximation: first applications with a skyrme energy functional. Phys. Rev. C 92, pp. 034303. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [17] (2019) Reference database for photon strength functions. The European Physical Journal A 55 (10). External Links: ISSN 1434-601X, Document Cited by: Figure 4.
- [18] (2014) Coupled-cluster computations of atomic nuclei. Rep. Prog. Phys. 77, pp. 096302. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [19] (2016) The in-medium similarity renormalization group: a novel ab initio method for nuclei. Physics Reports 621, pp. 165–222. External Links: ISSN 0370-1573 Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [20] (2011-06) Quasiparticle random-phase approximation with interactions from the similarity renormalization group. Phys. Rev. C 83, pp. 064317. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [21] (2016) Ab initio nuclear many-body perturbation calculations in the hartree-fock basis. Phys. Rev. C 94, pp. 014303. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [22] (2020) Accurate bulk properties of nuclei from to from potentials with isobars. Phys. Rev. C 102, pp. 054301. Cited by: Table 1, Table 2, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [23] (1993) Ground state correlations and charge transition densities. Physics Letters B 306 (3), pp. 197–200. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [24] (1990) RPA ground state correlations in nuclei. Physics Letters B 249 (3), pp. 377–380. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [25] (2023) On the dynamical kernels of fermionic equations of motion in strongly-correlated media. The European Physical Journal A 59 (12), pp. 291. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [26] (2016-04) Estimation of a 2p2h effect on gamow-teller transitions within the second tamm-dancoff approximation. Phys. Rev. C 93, pp. 044319. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [27] (2022-01) Converged ab initio calculations of heavy nuclei. Phys. Rev. C 105, pp. 014302. External Links: Document Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [28] (2023) NuHamil. Eur. Phys. J. A 59, pp. 150. External Links: Document Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [29] National nuclear data center database. External Links: Link Cited by: Figure 1, Figure 2.
- [30] (2006-07) Collective multipole excitations based on correlated realistic nucleon-nucleon interactions. Phys. Rev. C 74, pp. 014318. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [31] (2003-03) Quasiparticle random phase approximation based on the relativistic hartree-bogoliubov model. Phys. Rev. C 67, pp. 034312. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [32] (2007-01) Nuclear collective excitations using correlated realistic interactions: the role of explicit random-phase approximation correlations. Phys. Rev. C 75, pp. 014310. External Links: Document Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [33] (2010-02) Large-scale second random-phase approximation calculations with finite-range interactions. Phys. Rev. C 81, pp. 024317. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [34] (2019-05) Nuclear electromagnetic dipole response with the self-consistent green’s function formalism. Phys. Rev. C 99, pp. 054327. External Links: Document Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [35] (2004) The nuclear many-body problem. Springer Science & Business Media. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [36] (2018) Nuclear equation of state from ground and collective excited state properties of nuclei. Progress in Particle and Nuclear Physics 101, pp. 96. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [37] (1968-01) Equations-of-motion method and the extended shell model. Rev. Mod. Phys. 40, pp. 153–166. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [38] (1968-11) Methods for calculating ground-state correlations of vibrational nuclei. Phys. Rev. 175, pp. 1283–1292. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [39] (2010) Nuclear collective motion: models and theory. World Scientific. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [40] GitHub. Note: https://github.com/ragnarstroberg/imsrg Cited by: Table 1, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [41] (2021) Equation of motion method for strongly correlated fermi systems and extended rpa approaches. Physics Reports 929, pp. 1–84. External Links: Document Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [42] (1989) Reduced electric-octupole transition probabilities, B(E3; ), for even-even nuclides throughout the periodic table. Atomic Data and Nuclear Data Tables 42 (1), pp. 55–104. External Links: ISSN 0092-640X, Document Cited by: Table 3.
- [43] (2005) Relativistic hartree–bogoliubov theory: static and dynamic aspects of exotic nuclear structure. Physics Reports 409 (3), pp. 101. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
- [44] (2018-05) Chiral descriptions of nuclear multipole resonances within the random-phase approximation. Phys. Rev. C 97, pp. 054306. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.